Please help! Please explain with calculator inputs. Thank you.

You have just won a lawsuit. The settlement will pay either a single lump sum settlement today, or the
following cash flows (assume the following cash flows come at the end of the period):
Year 1: $10,000
Years 2-4: $ 5,000
Years 5 onward (forever): $ 3,000
You can presently earn 8% on your investments.
What is the least you should accept today as a lump sum settlement?
Answer #1

Let Cash Flows for Year n be CFn

Given, CF1 = 10000

CF2 = 5000

CF3 = 5000

CF4 = 5000

CF5 = 3000

CF6 ......

Amount to accept today as lump sum = Net Present Value of future Cash Flows

=> NPV = Σ CFn/(1+r)n, where r = 8%

=> NPV = 10000/(1+0.08) + 5000/(1+0.08)2 + 5000/(1+0.08)3 + 5000/(1+0.08)4 + 3000/(1+0.08)5 + ......

= 10000/(1+0.08) + 5000/(1+0.08)2 + 5000/(1+0.08)3 + 5000/(1+0.08)4 + (1/(1+0.08)5)3000/(1 - 1/(1+0.08))

= 10000/(1+0.08) + 5000/(1+0.08)2 + 5000/(1+0.08)3 + 5000/(1+0.08)4 + (1/(1+0.08)4)3000/0.08

= $48753.88